OpenStudy (anonymous):

If A is a rational number and B is a rational number then the product AB about must be? @CollateralDamage

OpenStudy (zzr0ck3r):

\(\dfrac{a}{b}(\dfrac{c}{d})=\dfrac{ac}{bd}\)

OpenStudy (anonymous):

is it a) rational B) imaginary C)an integer D)irrational

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

It should be rational... I think. I am pretty sure that's right since a true ratio (as in NOT one where it's like 3/1) should work out to be rational.